# Deep Learning-based MRI Reconstruction with Artificial Fourier Transform Network (AFTNet)

### v2

@@ Abstract
Deep complex-valued neural networks provide a powerful way to leverage complex number operations and representations and have succeeded in several phase-based applications. However, most previously published networks have not fully explored the impact of complex-valued networks in the frequency domain. Here, we introduce a unified complex-valued deep learning framework—Artificial Fourier Transform Network (AFTNet)—which combines domain-manifold learning and complex-valued neural networks. AFTNet can be readily used to solve image inverse problems in domain transformation, especially for accelerated magnetic resonance imaging (MRI) reconstruction and other applications. While conventional methods only accept magnitude images, the proposed method takes raw k-space data in the frequency domain as input, allowing a mapping between the k-space and image domains to be determined through cross-domain learning. We show that AFTNet achieves superior accelerated MRI reconstruction compared to existing approaches. Furthermore, our approach can be applied to various tasks, such as denoised magnetic resonance spectroscopy (MRS) reconstruction and datasets with various contrasts. The AFTNet presented here is a valuable preprocessing component for different preclinical studies and provides an innovative alternative for solving inverse problems in imaging and spectroscopy. The code is available at: https://github.com/yanting-yang/AFT-Net.
@@
journal: Medical Image Analysis
@@ 1 Introduction
The shift from real to complex coordinate space in deep neural networks has revealed the potential of complex numbers’ rich representational capacity, thereby advancing the development of complex-valued neural architectures[1, 2, 3]. A similar but inverse domain shift occurs in the preprocessing of magnetic resonance imaging (MRI), where raw data are acquired and stored in complex-valued k-space, with each pixel representing the spatial frequency information in two or three dimensions of an object. After data acquisition, the raw k-space data are transformed into images through the image reconstruction process, enabling interpretation by MR operators, physicians, radiologists, or data scientists. This reconstruction is an essential step in the preprocessing pipeline and lays the foundation for overall image quality. Proper image reconstruction methods can increase the signal-to-noise ratio (SNR) by removing thermal noise[4, 5], improve spatial inhomogeneities affected by point spread functions (PSFs), and correct unexpected signal artifacts[6].
@@
Theoretically, image data are reconstructed by domain transforms—for example, Fourier transforms for fully sampled Cartesian data[7]. However, under conditions of signal nonideality commonly encountered in clinical acquisitions, numerical and machine-learning methods are required. Human experts are involved in selecting task-related features and establishing models that accurately represent the mapping between the k-space and image domains[8], including methods such as Fast Fourier Transform (FFT)[9], noise prewhitening[6], digital signal processing (DSP), interpolation, and coil combination. Nevertheless, due to significant pathological variations and the potential for human oversight[10], a consistent and unbiased diagnosis cannot be guaranteed. In recent years, k-space, as a low-dimensional feature space, has been leveraged in deep neural networks to learn the manifold mapping of domain transforms in low signal-to-noise settings[11]. This image reconstruction process can be reformulated as a data-driven supervised learning task that determines the mapping between the k-space and image domains, demonstrating superior immunity to noise and reconstruction artifacts. The conventional discrete Fourier transform algorithm, which is mathematically derived and not based on learning, can be substituted by neural networks[12]. Thus, fundamental neural networks have been presented that avoid the difficulties of network structure finding and algorithm optimization. A similar approach is adopted in domain-transform manifold learning along the phase-encoding direction[13], where front-end convolutional layers, an intermediate global transform, and back-end convolutional layers are combined to perform data restoration in both k-space and image domains. Modern score-based diffusion models provide a powerful way to sample data from a conditional distribution given the measurements in the k-space domain, which can be used to solve inverse problems in imaging[14].
@@
While most deep learning-based MR image reconstruction algorithms have applied the concept of domain-transform learning that directly learns a low-dimensional joint manifold between the k-space and image domains, few works have fully leveraged complex-valued neural networks (CVNNs), which enable neural networks to implement learning-based frequency selection[15]. Some early works proposed CVNNs but mainly focused on solving the basics of learning[16, 17]. Extensive studies on complex-valued CNNs have been conducted in recent years. Generalizations of real-valued CNN models have been shown to be significantly less vulnerable to overfitting[2]. Mathematical arguments and implementations have also been discussed[18, 3], enabling the practical application of CVNNs. A combination of CVNNs and vision deep learning models (e.g., U-Net[19]) has recently emerged and is being exploited for MR image reconstruction[20, 21], demonstrating superior and accelerated reconstruction compared to real-valued neural networks. However, the main drawback of previous CVNN works is that the potential of leveraging CVNNs in domain manifold learning has not been fully investigated.
In this study, we combine domain-manifold learning with complex-valued neural networks to develop a unified end-to-end complex-valued image reconstruction approach for MRI. The framework we describe is the artificial Fourier transform network (AFTNet), as shown in Figure 1, which aims to facilitate the MR reconstruction with the presence of noise in k-space from the preprocessing workflow and incorporate data processing into deep learning frameworks. It consists of a trainable data-driven domain transformation block (AFT) that approximates the Fourier transform (FT) and complex-valued convolutional encoder-decoder networks to extract higher-level features in the k-space and image domains. Through its modular design, the proposed AFTNet can be extended to any dimension (e.g., 1D MR spectroscopy data). We use AFTNet to accelerate and denoise MR acquisitions. In accelerated MRI reconstruction, the k-space data are under-sampled in the data acquisition direction (the phase-encoding direction). We further apply the extended AFTNet to denoised MRS acquisition, where the free induction decay (FID) data are under-sampled by decreasing repetition numbers. Comprehensive experiments are conducted across various modalities, system field strengths, acceleration ratios, and noise levels.
@@
Figure 1: Schematics of general deep-learning MR imaging/spectroscopy reconstruction based on AFTNet. The generality of AFTNet is evaluated on the human normal-field MRI dataset (fastMRI[22]) and human normal-field MRS dataset (Big GABA[23]). The inputs of AFTNet can be 2D MRI k-space data or 1D MRS FID data. The outputs are reconstructed MR images or spectra. Different structures of AFTNet are developed by appending front-end and/or back-end convolutional networks to the AFT block. Here we show the human T2w 1.5T MRI and human 3T PRESS MRS reconstruction results, respectively. C: complex-valued and R: real-valued.
@@ 2 Background
2.1 Complex-valued neural network
The definition of the conventional real-valued neural network can be extended to the complex domain. We denote a complex operator as $\mathrm{W}=\mathrm{W}_{real}+i\mathrm{W}_{imag}$, where $\mathrm{W}_{real}$ and $\mathrm{W}_{imag}$ are real-valued operators. The input complex vector can the represented as $x=x_{real}+ix_{imag}$. The output of complex operator $\mathrm{W}$ acting on $x$ is derived following the vector dot product:
$\begin{split}y&=\mathrm{W}*x\\ &=(\mathrm{W}_{real}(x_{real})-\mathrm{W}_{imag}(x_{imag}))\\ &+i(\mathrm{W}_{imag}(x_{real})+\mathrm{W}_{real}(x_{imag})).\end{split}$ (1)
As the linear operator and convolution operator are distributive[3], we obtain:
$\begin{split}\mathbb{C}\mathrm{Linear}(z)&=(\mathrm{Linear}_{1}(\Re(z))-\mathrm{Linear}_{2}(\Im(z)))\\ &+i(\mathrm{Linear}_{2}(\Re(z))+\mathrm{Linear}_{1}(\Im(z)))\end{split}$ (2)
@@ 2.1 Complex-valued neural network
$\begin{split}\mathbb{C}\mathrm{Conv}(z)&=(\mathrm{Conv}_{1}(\Re(z))-\mathrm{Conv}_{2}(\Im(z)))\\ &+i(\mathrm{Conv}_{2}(\Re(z))+\mathrm{Conv}_{1}(\Im(z)))\end{split}$ (3)
where $z\in\mathbb{C}$ and we use subscripts $1$ and $2$ instead of $real$ and $imag$ to avoid misleading.
@@
Normalization is a common technique widely used in deep learning to accelerate training and reduce statistical covariance shift[24, 25, 26]. This is also mirrored in the complex-valued neural network, where we want to ensure that both the real and the imaginary parts have equal variance. Extending the normalization equation to matrix notation we have:
$\tilde{z}=V^{-\frac{1}{2}}(z-\mathbb{E}(z))$ (5)
and $V$ is the covariance matrix of real and imaginary parts of $z$
$V$ is a $2\times 2$ matrix, and the existence of the inverse square root is guaranteed by the addition of $\epsilon I$ (Tikhonov regularization). Therefore, the solution of the inverse square root can be expressed analytically as
where $s=\sqrt{AD-BC}$, $t=\sqrt{A+D+2s}$ and $d=st$.
The complex normalization is defined as
where $\gamma$ and $\beta$ are learnable parameters.
Considering the limitation of GPU RAM and large memory consumption of complex-valued networks, we use the group normalization[25] in our framework to avoid possible inaccurate batch statistics estimation caused by a small batch size. The group normalization in the complex domain can be represented as Algorithm 1.
@@ 3 Methods
Figure 2: Structure of 2-dimensional AFTNet. Components include the complex-valued AFT block, the complex-valued residual attention U-Net, the complex-valued residual block, and the complex-valued attention gate. All convolutional layers have a kernel size of 3, except those pointed out specifically. C: complex-valued. Red numbers indicate the number of channels produced by each layer.
@@ 3.1 Artificial Fourier Transform
where the $x$ and $y$ are two separable independent variables. Each dimension of 2D DFT can be modeled as a trainable neural network[12]. We further proposed that this idea could be naturally implemented with complex-valued neural networks shown in Figure 2.
@@
where $W_{real}$ and $W_{imag}$ are the real and the imaginary coefficients. Use matrix notation to represent real and imaginary parts of the DFT operation. We have:
@@
Compare Equation 13 with Equation 1, a multi-layer complex-valued neural network with linear activation function can represent 1D DFT with appropriate weights. We use $AFT_{N}$ to denote the complex-valued Fourier transform deep learning block on the input vector with $N$ elements. The Fourier transform of the input data with dimension $H\times W$ can be represented as
@@ 3.2 Network framework
The network structure and general workflow are shown in Figure 2 and Figure 3. The AFT block is composed of an MLP with three complex linear layers linked by two complex LeakReLU activation layers. We set the negative slope of LeakReLU to be $0.1$. As for conventional accelerated MRI acquisition, the undersampling is applied to the phase-encoding direction, we also only apply the AFT block to that direction. We apply our AFT to the multi-coil k-space data acquired directly from the scanner for the reconstruction task. The target is derived from the inverse fast Fourier transform on the input data. The AFT does not compress the coil channel so that the input and output shapes/sizes are the same. The network performance is evaluated within magnitude images obtained by the Fourier application and coil compression. For the accelerated reconstruction task, we combine our AFT with an entirely complex-valued residual attention U-Net (RAUNet)[20], which extracts local and global features in the k-space and/or image domain and forces the network to represent sparsely in those domains. Multiple network architectures are evaluated to verify the effectiveness of both AFT and RAUNet in different domains. We refer each of them to AFT, AFTNet-I, AFTNet-K, and AFTNet-KI, respectively as shown in Figure 3. We first train an AFT-only network to see if, without convolution layers, the AFT can remove artifacts and enhance quality. Then a network with the AFT followed by the RAUNet is trained to simulate a typical deep learning workflow where conventional numerical methods are used to preprocess the image, and CNNs are utilized to map the input domain to the target domain. We also evaluate the network with RAUNet first implemented directly on the k-space domain. Given that each position in k-space contains the information of the whole image, CNNs implemented in k-space can leverage the complete information of all space, even if they have a fixed field of view. Finally, a RAUNet-AFT-RAUNet structure is evaluated with the first RAUNet extracting k-space domain features and the second RAUNet extracting image domain features.
The architecture of the RAUNet presented in Figure 2 is generally based on the residual attention U-Net[19] but with all the real-valued components replaced by complex-valued components as shown in Figure 2, including complex-valued convolutional layers and complex-valued ReLU layers introduced in Section 2.1. We further optimize the network for smaller batch sizes by replacing batch normalization with group normalization as illustrated in Algorithm 1. Other complex components are implemented in the same way. For example, the complex transposed convolution operator can be mirrored from Equation 3, complex sigmoid is applied like complex ReLU, and complex max pooling is almost the same as the real-valued version except that indices are inferred from absolute values.
3.3 Implementation details
We construct a batch size of 1 and optimize the network using the ADAM[27] optimizer. We use a learning rate scheduler with cosine annealing[28]. The initial learning rate is $10^{-3}$ and the final learning rate is $10^{-5}$. All methods are trained for 50 epochs. All experiments are done using PyTorch 2.2.2 and a Quadro RTX 6000 GPU.
In the context of image reconstruction and processing, the impact of the loss function is vital if the final results are to be evaluated by human observers. One common and safe choice is $\ell_{2}$ loss which works under the assumption of white Gaussian noise. For training AFT for MRI reconstruction, the loss value is determined in the frequency domain as
@@ 3.3 Implementation details
Data consistency is applied after each RACUNet and AFT block so that the known positions of output data are replaced by the original data samples to get better fidelity[31].

### v3

@@ Abstract
Deep complex-valued neural networks (CVNNs) provide a powerful way to leverage complex number operations and representations and have succeeded in several phase-based applications. However, previous networks have not fully explored the impact of complex-valued networks in the frequency domain. Here, we introduce a unified complex-valued deep learning framework—Artificial Fourier Transform Network (AFTNet)—which combines domain-manifold learning and CVNNs. AFTNet can be readily used to solve image inverse problems in domain transformation, especially for accelerated magnetic resonance imaging (MRI) reconstruction and other applications. While conventional methods typically utilize magnitude images or treat the real and imaginary components of k-space data as separate channels, our approach directly processes raw k-space data in the frequency domain, utilizing complex-valued operations. This allows for a mapping between the frequency (k-space) and image domain to be determined through cross-domain learning. We show that AFTNet achieves superior accelerated MRI reconstruction compared to existing approaches. Furthermore, our approach can be applied to various tasks, such as denoised magnetic resonance spectroscopy (MRS) reconstruction and datasets with various contrasts. The AFTNet presented here is a valuable preprocessing component for different preclinical studies and provides an innovative alternative for solving inverse problems in imaging and spectroscopy. The code is available at: https://github.com/yanting-yang/AFT-Net.
@@
journal: Computers in Biology and Medicine
@@ 1 Introduction
The shift from real to complex coordinate space in deep neural networks has unveiled the potential of complex numbers’ rich representational capacity, thereby spurring the development of complex-valued neural architectures [1, 2, 3]. A corresponding, though inverse, domain shift occurs in the preprocessing of magnetic resonance imaging (MRI), where raw data are acquired and stored in complex-valued k-space, with each pixel encoding spatial frequency information across two or three dimensions. Following acquisition, the raw k-space data are transformed into images via a reconstruction process that enables interpretation by MR operators, physicians, radiologists, or data scientists. This crucial reconstruction step underpins overall image quality. Proper image reconstruction methods can increase the signal-to-noise ratio (SNR) by suppressing thermal noise [4, 5], improve spatial inhomogeneities affected by point spread functions (PSFs), and correct unexpected signal artifacts [6].
@@
Theoretically, image reconstruction is performed using domain transforms, for example Fourier transforms for fully sampled Cartesian data [7]. However, in clinical settings where signal non-idealities are prevalent, numerical methods and machine learning approaches become indispensable. Traditionally, human experts select task-related features and develop models to capture the mapping between k-space and image domains [8]. However, due to significant pathological variations and the risk of human oversight [9], a consistent and unbiased diagnosis cannot be guaranteed. In recent years, k-space, as a low-dimensional feature space, has been leveraged in deep neural networks to learn the manifold mapping of domain transforms in low signal-to-noise settings [10]. This image reconstruction process can be reformulated as a data-driven supervised learning task that determines the mapping between the k-space and the image domains, demonstrating superior immunity to noise and reconstruction artifacts. The conventional discrete Fourier transform algorithm, which is mathematically derived and is not based on learning, can be substituted by neural networks [11], as neural networks can learn complex mappings and priors from data, enabling them to recover missing information and reduce artifacts more effectively. This is particularly important in highly undersampled scenarios, where traditional DFT-based methods struggle to achieve high-quality reconstructions. Thus, fundamental neural networks have been presented that avoid the difficulties of finding the network structure and optimizing the algorithm. A similar approach is adopted in domain transform manifold learning along the phase-encoding direction [12], where the front-end convolutional layers, an intermediate global transform, and the back-end convolutional layers are combined to perform data restoration in the k-space and image domains. Modern score-based diffusion models provide a powerful way to sample data from a conditional distribution given the measurements in the k-space domain, which can be used to solve inverse problems in imaging [13].
@@
Although most deep learning-based MR image reconstruction algorithms have applied the concept of domain-transform learning that directly learns a low-dimensional joint manifold between the k-space and image domains, including end-to-end [14] and the unrolled [15] network, few works have fully leveraged complex-valued neural networks (CVNNs), which allow neural networks to implement learning-based frequency selection [16]. Some early works proposed CVNNs but focused mainly on solving the basics of learning [17, 18]. In recent years, extensive studies have been conducted on complex-valued CNNs. Generalizations of real-valued CNN models have been shown to be significantly less vulnerable to overfitting [2]. Mathematical arguments and implementations have also been discussed [19, 3], enabling the practical application of CVNNs. A combination of CVNNs and vision deep learning models (e.g., UNet [20]) has recently emerged and is being exploited for MR image reconstruction [21, 22], demonstrating superior and accelerated reconstruction compared to real-valued neural networks. However, the main drawback of previous CVNN works is that the potential of leveraging CVNNs in domain manifold learning has not been fully investigated.
To fully exploit the rich information contained in complex-valued MR data and overcome the limitations of traditional reconstruction methods, we propose (1) Artificial Fourier Transform (AFT) block which replaces conventional DFT with a learnable module and can adapt to noise and signal non-idealities, allowing the network to learn complex mappings between the k-space and image domains; (2) Artificial Fourier Transform Network (AFTNet), shown in Figure 1, which combines domain-manifold learning with CVNNs by integrating AFT with complex-valued convolutional encoder-decoder networks to facilitate MR reconstruction. AFTNet directly process the full complex information (both magnitude and phase) inherent in k-space, avoiding the common pitfall of separating or discarding phase information. By operating directly on raw k-space data, AFTNet minimizes information loss and learns subtle frequency domain features critical for accurate image reconstruction. The CVNNs in AFTNet process global features in the frequency (k-space) domain and local details in the image domain, leading to improved artifact removal and reconstruction quality, particularly in highly undersampled scenarios. Through its modular design, AFTNet can be extended to any dimension (e.g., 1D MR spectroscopy data). We conduct extensive experiments in various modalities, system field strengths, acceleration ratios, and noise levels on the MRI reconstruction, accelerated reconstruction and MRS denoised reconstruction, achieving superior performance in all scenarios.
@@
Figure 1: Schematics of general deep-learning MR imaging/spectroscopy reconstruction based on AFTNet. The inputs of AFTNet can be 2D MRI k-space data or 1D MRS FID data. The outputs are reconstructed MR images or spectra. Different structures of AFTNet are developed by appending front-end and/or back-end convolutional networks to the AFT block. Here we show the T2w 1.5T MRI and 3T PRESS MRS reconstruction results, respectively. C: complex-valued and R: real-valued.
@@ 2 Background
2.1 Complex-valued neural networks
The definition of a conventional real-valued neural network can be extended to the complex domain. We denote a complex operator as $\mathrm{W}(\cdot)=\mathrm{W}_{real}(\cdot)+i\mathrm{W}_{imag}(\cdot)$, where $\mathrm{W}_{real}$ and $\mathrm{W}_{imag}$ are real-valued operators. The input complex vector can be represented as $z=\Re(z)+i\Im(z)$. The output of complex operator $\mathrm{W}$ acting on $z$ is derived following the vector dot product:
$\begin{split}y&=\mathrm{W}\cdot z\\ &=\left[\mathrm{W}_{\text{real}}(\Re(z))-\mathrm{W}_{\text{imag}}(\Im(z))\right]\\ &\quad+i\left[\mathrm{W}_{\text{imag}}(\Re(z))+\mathrm{W}_{\text{real}}(\Im(z))\right].\end{split}$ (1)
As the linear operator and convolution operator are distributive [3], we can directly replace $\mathrm{W}(\cdot)$ with $\mathrm{Linear}(\cdot)$ or $\mathrm{Conv}(\cdot)$ to obtain the complex-valued version:
$\begin{split}\mathbb{C}\mathrm{Linear}(z)&=\left[\mathrm{Linear}_{1}(\Re(z))-\mathrm{Linear}_{2}(\Im(z))\right]\\ &\quad+i\left[\mathrm{Linear}_{2}(\Re(z))+\mathrm{Linear}_{1}(\Im(z))\right]\end{split}$ (2)
@@ 2.1 Complex-valued neural networks
$\begin{split}\mathbb{C}\mathrm{Conv}(z)&=\left[\mathrm{Conv}_{1}(\Re(z))-\mathrm{Conv}_{2}(\Im(z))\right]\\ &\quad+i\left[\mathrm{Conv}_{2}(\Re(z))+\mathrm{Conv}_{1}(\Im(z))\right]\end{split}$ (3)
where we use subscripts $1$ and $2$ instead of $real$ and $imag$ to avoid misleading.
@@
Normalization is a common technique widely used in deep learning to accelerate training and reduce the statistical covariance shift [23, 24, 25]. This is also mirrored in the CVNNs, where we want to ensure that both the real and the imaginary parts have equal variance. Extending the normalization equation to matrix notation we have:
$\tilde{z}=V^{-\frac{1}{2}}(z-\mathbb{E}(z))$ (5)
and $V$ is the covariance matrix of real and imaginary parts of $z$
$V$ is a $2\times 2$ matrix, and the existence of the inverse square root is guaranteed by adding $\epsilon I$ (Tikhonov regularization). Therefore, the solution of the inverse square root can be expressed analytically as
where $s=\sqrt{AD-BC}$, $t=\sqrt{A+D+2s}$ and $d=st$.
The complex normalization is defined as
where $\gamma$ and $\beta$ are learnable parameters.
Considering the limitation of GPU RAM and the large memory consumption of complex-valued networks, we use group normalization [24] in our framework to avoid the possible inaccurate estimation of batch statistics caused by a small batch size. The group normalization in the complex domain can be represented as Algorithm 1.
@@ 3.1 Artificial Fourier Transform
Figure 2: Structure of 2-dimensional AFTNet. Components include the complex-valued AFT block, the complex-valued residual attention UNet, the complex-valued residual block, and the complex-valued attention gate. All convolutional layers have a kernel size of 3, except those pointed out specifically. C: complex-valued. Red numbers indicate the number of channels produced by each layer.
@@
where the $x$ and $y$ are two separable independent variables. Each dimension of 2D DFT can be modeled as a trainable neural network [11]. We further proposed that this idea could be naturally implemented with CVNNs shown in Figure 2.
@@
where $W_{real}$ and $W_{imag}$ are the real and the imaginary coefficients. Use matrix notation to represent the real and imaginary parts of the DFT operation. We have the following:
@@
Compared Equation 13 with Equation 1, a multilayer CVNN with linear activation function can represent 1D DFT with appropriate weights. We use $AFT_{N}$ to denote the complex-valued Fourier transform deep learning block on the input vector with $N$ elements. The Fourier transform of the input data with dimension $H\times W$ can be represented as
@@ 3.2 Network framework
The network structure and general workflow are shown in Figure 2 and Figure 3. The AFT block is composed of an MLP with three complex linear layers linked by two complex LeakReLU activation layers. As for conventional accelerated MRI acquisition, the undersampling is applied to the phase-encoding direction, and we also only apply the AFT block to that direction. For the accelerated reconstruction task, we combine our AFT with an entirely complex-valued UNet (CUNet) [21] with residual attention, which extracts local and global features in the k-space and/or image domain. Multiple network architectures are evaluated to verify the effectiveness of both AFT and CUNet in different domains. We refer each of them to AFT, AFTNet-I, AFTNet-K, and AFTNet-KI, respectively as shown in Figure 3.
The architecture of the CUNet presented in Figure 2 is generally based on the residual attention UNet [20] but with all the components of real value replaced by complex value components as shown in Figure 2, including complex-valued convolutional layers and complex-valued ReLU layers introduced in Section 2.1. We further optimize the network for smaller batch sizes by replacing batch normalization with group normalization as illustrated in Algorithm 1. Other complex components are implemented in the same way. For example, the complex transposed convolution operator can be mirrored from Equation 3, the complex sigmoid is applied like the complex ReLU, and the complex max pooling is almost the same as the real-valued version except that the indices are inferred from absolute values.
3.3 Reconstruction Loss
In the context of image reconstruction and processing, the impact of the loss function is vital if human observers are to evaluate the final results. One common and safe choice is $\ell_{2}$ loss which works under the assumption of Gaussian white noise. For training AFT for MRI reconstruction, the loss value is determined in the frequency domain as
@@ 3.3 Reconstruction Loss
Data consistency is applied after each RACUNet and AFT block so that the known positions of the output data are replaced by the original data samples to obtain better fidelity [28].
4 Experiments
4.1 Overview
In the MRI reconstruction experiment, we apply our AFT to the multicoil k-space data acquired directly from the scanner for the reconstruction task. The target is derived from the fast inverse Fourier transform on the input data. The AFT does not compress the coil channel so that the input and output shapes/sizes are the same. Network performance is evaluated within magnitude images obtained by the Fourier application and coil compression, which are crucial for radiologists in clinical applications. We also recognize some clinical settings that rely on the phase image, so we also provide evaluation on the phase images.
For the MRI accelerated reconstruction experiment, we first trained AFT, CUNet (DFTNet-I) and DFTNet-K models. Then, we replace the DFT with AFT in the DFTNet to train the AFTNet (I and K) models. At last, we concatenate the pretrained K model with I model to train the KI model. To be more specific, we train an AFT-only network to see if, without convolution layers, the AFT can remove artifacts and enhance quality. Then a network with the AFT followed by the CUNet is trained to simulate a typical deep learning workflow where conventional numerical methods are used to preprocess the image and CNNs are utilized to map the input domain to the target domain. We also evaluate the network with CUNet first implemented directly in the k-space domain. Given that each position in k-space contains the information of the whole image, CNNs implemented in k-space can leverage the complete information of all spaces, even if they have a fixed field of view. Finally, a CUNet-AFT-CUNet structure is evaluated with the first CUNet that extracts k-space domain features and the second CUNet extracting image domain features.
We also extended AFTNet to 1D scenario for spectral reconstruction to evaluate its generality. In such case, we use the AFTNet-I model and train it from scratch.